Q5: Bar Chart 

The number of bed-sheets manufactured by a factory during five consecutive weeks is given below. 

Week First Second Third Fourth Fifth 

Number of Bed-sheets 600 850 700 300 900 

Draw the bar graph representing the above data. 

Q6: Bar Chart 

The number of absentees in class VIII was recorded in a particular week. Represent this data  on the bar graph 

Days Mon. Tues. Wed. Thurs. Fri. Sat. 

Number of Absentees 130 120 135 130 150 80 

a. On which day the maximum and minimum students were absent? 

b. How many students were absent on Wednesday and Friday? 

Q7: Pie Chart 

The following table shows the mode of transport used by 400 students of a school. Represent  the following information on the pie chart. (Show the steps of construction of pie graph for  the given data along with the calculation)

Mode of Transport Bus Bicycle On foot By car 

No. of Students 200 100 80 20


1
Expert's answer
2021-02-24T14:25:57-0500

Q5



Q6



a. The maximum number of students were absent on Friday (150 students) and the minimum number of students were absent on Saturday (80 students).

b. On Wednesday and Friday 135 + 150 = 285 students were absent.


Q7



Total number of students = 400

Bus: "\frac{200}{400}\cdot 100\% = 50\%"


Bicycle: "\frac{100}{400}\cdot 100\% = 25\%"


On foot: "\frac{80}{400}\cdot 100\% = 20\%"


By car: "\frac{20}{400}\cdot 100\% = 5\%"